# Scalable multiphoton generation from cavity-synchronized single-photon sources

### v2

@@ II.3 Synchronization dynamics and parameter regime
Our goal is to realize the synchronized excitation and emission of each qubit, so that they act as nearly independent SPSs. The first stage of operation involves exciting each two-level systems from the ground state $|g\rangle_{j}$ to the excited state $|e\rangle_{j}$ via a fast cavity-mediated $\pi$-pulse of duration $T\sim\pi/(2\Omega_{\rm cm}^{j})$. We then expect that all emitters will produce synchronized photons, one on each of the antennas, on a timescale $t\gg 1/\gamma_{j}$. To ensure that this procedure efficiently generates $N$ nearly indistinguishable and independent single-photons, the system parameters must satisfy the conditions,
@@
First, to achieve synchronization and indistinguishability of the photon emissions, the system parameters should be as homogeneous as possible, and specially the qubit frequencies $\omega_{q}\approx\omega_{q}^{j}$, couplings $g\approx g_{j}$, and antenna decays $\gamma\approx\gamma_{j}$. In the remainder of the paper we will thus assume that all system parameters are homogeneous, except in Sec. V, where we analyze inhomogeneities and disorder.
@@
Figure 2: Synchronization of two deterministic SPSs via a cavity. (a) Short time dynamics: The cavity pulse is switched on $(f(t)=1)$ and a large coherent state with $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|^{2}\sim 10^{4}$ photons is quickly created in the cavity with fast residual oscillations due to non-RWA terms [cf. red solid and black dashed lines]. This induces a cavity-mediated driving $\Omega_{\rm cm}=|\alpha_{\rm ss}|g$, which resonantly excites both qubits $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle$ synchronously [cf. blue dash-dotted line]. After a $\pi$-pulse time $T\sim\pi/(2\Omega_{\rm cm})$, the cavity pulse is switched off $(f(t)=0)$, $\langle a^{\dagger}a\rangle$ decays in a fast timescale $\sim 1/\kappa$, and both qubits are left in their excited states $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$. (b) Long time dynamics: Both qubits spontaneously decay in a timescale $t\sim 15/\gamma$, and each of them generates a single-photon with probability ${\cal P}_{1}\approx 0.99$ on its own antenna [cf.  green dashed line]. Since these spontaneous emission processes are much faster than the timescale for the cavity-mediated interactions ($J_{\rm cm}/\gamma\sim 10^{-5}$) and collective decay ($\gamma_{\rm cm}/\gamma\sim 10^{-4}$), both single-photons are nearly perfectly synchronized and independent as manifested by the two-photon probability satisfying $P_{2}\approx({\cal P}_{1})^{2}\approx 0.98$ [cf. yellow dotted line]. (c)-(d) Deviations $D(x)=|x_{\rm exact}-x_{\rm eff}|$ between the exact (6) and effective (12) models for the qubit occupations $x=\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle$, and photon generation probabilities $x={\cal P}_{1},P_{2}$ as a function of time. We observe great agreement and no heating effects despite the large cavity driving strength $\Omega_{0}=4\omega_{d}\gg\kappa,|\Delta|$. These and the rest of the parameters correspond to a state-of-the-art circuit-QED implementation shown in row C of Table 1 and discussed in Sec. III.1.
$\frac{\gamma}{2\pi}[{\rm MHz}]$ $\frac{\gamma_{\rm loss}}{2\pi}[{\rm kHz}]$ $\frac{\gamma_{\phi}}{2\pi}[{\rm kHz}]$ $\gamma/\Omega_{\rm cm}$ $\gamma_{\phi}/\gamma$ $\gamma_{\rm loss}/\gamma$ $\gamma_{\rm cm}/\gamma$ $J_{\rm cm}/\gamma$ ${\cal P}_{1}$ $P_{2}$ $g^{(2)}_{\rm HBT}[0]$ $g^{(2)}_{\rm HOM}[0]$
A 0.4 0 0 $1.0\cdot 10^{-2}$ 0 0 $9.4\cdot 10^{-4}$ $1.2\cdot 10^{-4}$ 0.997 0.994 $4.4\cdot 10^{-3}$ $2.0\cdot 10^{-5}$
B 0.4 1 1 $1.0\cdot 10^{-2}$ $2.5\cdot 10^{-3}$ $2.5\cdot 10^{-3}$ $9.4\cdot 10^{-4}$ $1.2\cdot 10^{-4}$ 0.994 0.989 $4.4\cdot 10^{-3}$ $2.5\cdot 10^{-3}$
C 1.0 5 10 $2.6\cdot 10^{-2}$ $1.0\cdot 10^{-2}$ $5.0\cdot 10^{-3}$ $3.8\cdot 10^{-4}$ $4.7\cdot 10^{-5}$ 0.989 0.979 $1.0\cdot 10^{-2}$ $9.7\cdot 10^{-3}$
D 3.0 50 100 $7.7\cdot 10^{-2}$ $3.3\cdot 10^{-2}$ $1.7\cdot 10^{-2}$ $1.3\cdot 10^{-4}$ $1.6\cdot 10^{-5}$ 0.968 0.936 $3.0\cdot 10^{-2}$ $3.1\cdot 10^{-2}$
@@
Table 1: Parameters and figures of merit for a circuit-QED implementation of two synchronized SPSs. We study four parameter sets: A, B, C, and D, which show increasing antenna coupling $\gamma$ and decoherence rates $\gamma_{\rm loss}$, $\gamma_{\phi}$. The rest of the parameters are common to all sets and read $g/2\pi=0.2{\rm MHz}$, $\kappa/2\pi=400{\rm MHz}$, $\Delta/2\pi=50{\rm MHz}$, and $\Omega_{0}/2\pi=40{\rm GHz}$, which leads to $|\alpha_{\rm ss}|^{2}\approx 3.8\cdot 10^{4}$ photons in the cavity on average. The detuning of the qubits is chosen to compensate for the cavity-induced Lamb-shift $\delta_{q}=\delta_{\rm cm}$. Depending on the decoherence rates of sets A-D, the inequalities in Eq. (18) are satisfied differently, as shown in columns 5th to 9th, and this influences the performance of the four figures of merit: single-photon generation probability ${\cal P}_{1}$ [cf. Sec. III.2], synchronized two-photon generation efficiency or probability $P_{2}\approx({\cal P}_{1})^{2}$ [cf. Sec. III.3], multi-photon contamination $g^{(2)}_{\rm HBT}[0]$ [cf. Sec. III.4], and photon distinguishability $g^{(2)}_{\rm HOM}[0]$ [cf. Sec. III.5]. All these quantities are calculated using the full master equation (6), the methods in Appendices B.1 and C, and evolving up to a final time $t=15/\gamma$.
@@
Fig. 2(a)-(b) illustrates the two processes in the cavity-mediated synchronization method, using two synchronized SPSs with parameters that satisfy Eq. (18). Fig. 2(a) displays the short time dynamics: the creation of a large coherent state on the cavity $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|^{2}\sim 10^{4}$ [cf. red solid line], and the cavity-mediated simultaneous excitation of the qubits $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$ [cf. blue dash-dotted line]. Fig. 2(b) shows the long time dynamics, in which the photons are generated. Each emitter decays almost independently in a timescale $t\sim 15/\gamma$, depositing a photon into its own antenna [cf. blue dash-dotted line]. To quantify the efficiency of these emission processes, we display the single-photon and synchronized two-photon generation probabilities, ${\cal P}_{1}$ and $P_{2}$, calculated using the definitions and methods in Secs. III.2-III.3. For non-optimized parameters satisfying Eq. (18), we reach ${\cal P}_{1}\approx 0.99$ [cf. green dashed line] and $P_{2}\approx({\cal P}_{1})^{2}\approx 0.98$ [cf. yellow dotted line]. This demonstrates that both emitted photons are highly independent between each other $S_{2}\approx 1$ in Eq. (2)] and that cavity-induced correlations are indeed negligible during photon emission processes ($J_{\rm cm}/\gamma\sim 10^{-5}$ and $\gamma_{\rm cm}/\gamma\sim 10^{-4}$).
@@
We performed the numerical simulations in Figs. 2(a)-(b) using both the full model in Eqs. (4)-(6) and the effective model in Eqs. (12)-(17). They show an excellent agreement with deviations $D(x)=|x_{\rm exact}-x_{\rm eff}|<10^{-2}$ in the qubit and photon counting observables [cf. Fig. 2(c)-(d)], despite the large cavity population of $\langle a^{\dagger}a\rangle\sim 10^{4}\gg 1$. This demonstrates that the strong cavity driving $\Omega_{0}\gg\kappa,|\Delta|$ does not introduce heating in the system, but only creates a large coherent state in the cavity with some fast oscillations around the mean value $|\alpha_{\rm ss}|^{2}\gg 1$ [cf. solid red and dashed black lines in Fig. 2(a)]. Since this fast oscillations have no detrimental impact in the functionality of our multiphoton generation scheme, in the remainder of the paper we use the simpler effective model (12) which uses the mean steady value (9) for describing the effect of the cavity field (see App. A for more details on the approximations involved in the model).
@@ III.1 State-of-the-art parameters
In Fig. 1(b) we sketch a possible setup to implement the synchronized multiphoton emitter using superconducting circuits. For the cavity mode $a$, we consider a superconducting transmission line resonator with high decay rate $\kappa/2\pi=400{\rm MHz}$, and resonance frequency $\omega_{c}$ on the order of a few GHz. These superconducting resonators support strong drives and can hold a large number of photons in steady state $|\alpha_{\rm ss}|^{2}\gg 1$ Pietikäinen et al. (2017); Jenkins et al. (2014); Macrì et al. (2020), as it is required in our scheme. For this, we consider a driving with strength $\Omega_{0}/2\pi=40{\rm GHz}$, a detuning $\Delta/2\pi=50{\rm MHz}$, and an average number of photons $|\alpha_{\rm ss}|^{2}\approx 3.8\cdot 10^{4}$, well below the values $10^{5}-10^{9}$ photons observed in similar experiments Pietikäinen et al. (2017); Jenkins et al. (2014).
@@
As two-level emitters, we consider flux-tunable transmon qubits DiCarlo et al. (2009); Barends et al. (2013); Pechal et al. (2014); Wang et al. (2020) weakly coupled to the resonator mode with strength $g/2\pi=0.2{\rm MHz}$. The cavity mediated driving amplitude is $\Omega_{\rm cm}/2\pi=|\alpha_{\rm ss}|g/2\pi\approx 39$MHz. This is strong but still smaller than typical transmon anharmonicities $U/2\pi\sim 300-400{\rm MHz}$ Pechal et al. (2014); McKay et al. (2019); Zhou et al. (2020), thus preventing leakage to higher excitation states. The frequency of the qubits—which will be on the few GHz range—can be fine-tuned by sending a DC current through the antenna that is coupled to each qubit [cf. Fig. 1(b)]. This is required to bring the qubits in resonance with the external drive, and to compensate the cavity-induced Lamb-shifts $\delta_{q}=\delta_{\rm cm}$ [See Sec. V for small deviations of this condition].
The antenna of each qubit is not only used for frequency calibration, but its main purpose is to collect the photons emitted by each qubit [cf. Fig. 1(b)]. The coupling rate between the qubit and antenna $\gamma$ can be chosen depending on the estimated decoherence parameters $\gamma_{\rm loss}$ and $\gamma_{\phi}$, so that the photons are efficiently ($\gamma\gg\gamma_{\rm loss}$) and coherently ($\gamma\gg\gamma_{\phi}$) collected. In our analysis, we consider four different combinations of the parameter sets ($\gamma,\gamma_{\rm loss}$$\gamma_{\phi}$), shown in Table 1. Parameter set D corresponds to the typical relaxation $\gamma_{\rm loss}/2\pi\sim 50$kHz and dephasing rates $\gamma_{\phi}/2\pi\sim 100$kHz found in microwave quantum photonics experiments Pechal et al. (2014); Zhou et al. (2020); Wang et al. (2020). For this case we choose a relatively large decay rate of $\gamma/2\pi=3$MHz so that the emission time is faster than the decoherence timescales, but at the same time it is still slower than the effective driving $\Omega_{\rm cm}/2\pi\sim 39$MHz [cf. column D in Table 1]. In addition, parameter set C corresponds to decoherence rates reported in quantum computing experiments Barends et al. (2013); Caldwell et al. (2018); McKay et al. (2019); Andersen et al. (2020), which have reached relaxation and dephasing rates as low as $\gamma_{\rm loss}/2\pi\sim 5$kHz and $\gamma_{\phi}/2\pi\sim 10$kHz, respectively. For these parameters we can choose a slightly lower coupling to the antenna of $\gamma/2\pi\sim 1$MHz, and thus improve the ratio $\gamma/\Omega_{\rm cm}\sim 2.6\cdot 10^{-2}$ [cf. column C in Table 1]. Finally, set B corresponds to near-future parameters and set A to the ideal case of no decoherence, which we use as a reference to quantify the intrinsic limitations of the synchronization method. In the remainder of the paper, we use these four parameter sets A-D to characterize the performance and scaling of the synchronized SPSs for this circuit-QED implementation.
@@
Figure 3: Few-photon statistics ${\cal P}_{n}$ and two-photon dependence $D_{2}$ as a function of various system parameters such as (a) cavity occupation $|\alpha_{\rm ss}|^{2}$, (b) qubit-cavity coupling $g/\kappa$, (c) qubit-antenna decay $\gamma/\Omega_{\rm cm}$, (d) qubit-drive detuning $(\delta-\delta_{\rm cm})/\Omega_{\rm cm}$, (e) dephasing $\gamma_{\phi}/\gamma$, and (f) relaxation loss $\gamma_{\rm loss}/\gamma$. The probabilities ${\cal P}_{0}$ (green), $1-{\cal P}_{1}$ (blue), and ${\cal P}_{2}$ (brown) are plotted using the left vertical axis, whereas the two-photon dependence $D_{2}$ (red), corresponds the right vertical axis. Solid lines are computed using the effective master equation in Eq. (12) and the discrete symbols using the full master equation in Eq. (6), which show excellent agreement. Parameters are given by set C of Table 1, except for the quantity that is changed on each subplot.

### v3

@@ II.3 Synchronization dynamics and parameter regime
Our goal is to realize the synchronized excitation and emission of each qubit so that they act as nearly independent SPSs. The first stage of operation involves exciting each two-level system from the ground state $|g\rangle_{j}$ to the excited state $|e\rangle_{j}$ via a fast cavity-mediated $\pi$-pulse of duration $T\sim\pi/(2\Omega_{\rm cm}^{j})$. We then expect that all emitters will produce synchronized photons, one on each of the antennas, on a timescale $t\gg 1/\gamma_{j}$. To ensure that this procedure efficiently generates $N$ nearly indistinguishable and independent single-photons, the system parameters must satisfy the conditions,
@@
First, to achieve synchronization and indistinguishability of the photon emissions, the system parameters should be as homogeneous as possible, especially the qubit frequencies $\omega_{q}\approx\omega_{q}^{j}$, couplings $g\approx g_{j}$, and antenna decays $\gamma\approx\gamma_{j}$. In the remainder of the paper, we will thus assume that all system parameters are homogeneous, except in Sec. V, where we analyze inhomogeneities and disorder.
@@
Figure 2: Synchronization of two deterministic SPSs via a cavity. (a) Short time dynamics: The cavity pulse is switched on $(f(t)=1)$ and a large coherent state with $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|^{2}\sim{\color[rgb]{0,0,0}10^{2}}$ photons is quickly created in the cavity with fast oscillations due to the large detuning $\Delta$ and the non-RWA terms [cf. red solid and black dashed lines]. This induces an effective cavity-mediated driving $\Omega_{\rm cm}=|\alpha_{\rm ss}|g$, which resonantly excites both qubits $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle$ synchronously [cf. blue dash-dotted line]. After a $\pi$-pulse time $T\sim\pi/(2\Omega_{\rm cm})$, the cavity pulse is switched off $(f(t)=0)$, $\langle a^{\dagger}a\rangle$ decays in a fast timescale $\sim 1/\kappa$, and both qubits are left in their excited states $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$. (b) Long-time dynamics: Both qubits spontaneously decay in a timescale $t\sim 15/\gamma$, and each of them generates a single-photon with probability ${\cal P}_{1}\approx 0.99$ on its antenna [cf.  green dashed line]. Since these spontaneous emission processes are much faster than the timescale for the cavity-mediated interactions ($J_{\rm cm}/\gamma\sim{\color[rgb]{0,0,0}10^{-2}}$) and collective decay ($\gamma_{\rm cm}/\gamma\sim 10^{-4}$), both single-photons are nearly perfectly synchronized and independent as manifested by the two-photon probability satisfying $P_{2}\approx({\cal P}_{1})^{2}\approx 0.98$ [cf. yellow dotted line]. (c)-(d) Deviations $D(x)=|x_{\rm exact}-x_{\rm eff}|$ between the exact (7) and effective (13) models for the qubit occupations $x=\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle$, and photon generation probabilities $x={\cal P}_{1},P_{2}$ as a function of time. We observe great agreement and no heating effects despite the large cavity driving strength $\Omega_{0}=4\omega_{d}\gg\kappa,|\Delta|$. These and the rest of the parameters correspond to a state-of-the-art circuit-QED implementation shown in row A of Table 1 and discussed in Sec. III.1.
$g/2\pi$ $\kappa/2\pi$ $\Delta/2\pi$ $|\alpha_{\rm ss}|^{2}$ $\gamma/2\pi$ $\gamma_{\rm loss}/2\pi$ $\gamma_{\phi}/2\pi$ $\gamma/\Omega_{\rm cm}$ $\gamma_{\rm cm}/\gamma$ $J_{\rm cm}/\gamma$ ${\cal P}_{1}$ $P_{2}$ $g^{(2)}_{\rm HBT}[0]$ $g^{(2)}_{\rm HOM}[0]$
$\rm MHz$}] $\rm MHz$}] $\rm MHz$}] $\rm kHz$}] $\rm kHz$}] $\rm kHz$}]
A 4.0 50 1400 $10^{2}$ 966 5 10 $2.4\cdot 10^{-2}$ $4.2\cdot 10^{-4}$ $1.2\cdot 10^{-2}$ 0.989 0.979 $1.4\cdot 10^{-2}$ $1.4\cdot 10^{-2}$
B 4.0 50 1400 $10^{2}$ 3000 10 20 $7.5\cdot 10^{-2}$ $1.4\cdot 10^{-4}$ $3.8\cdot 10^{-3}$ 0.982 0.964 $2.8\cdot 10^{-2}$ $7.0\cdot 10^{-3}$
C 0.2 400 50 $4\cdot 10^{4}$ 400 1 1 $1.0\cdot 10^{-2}$ $9.4\cdot 10^{-4}$ $1.2\cdot 10^{-4}$ 0.994 0.989 $4.4\cdot 10^{-3}$ $2.5\cdot 10^{-3}$
D 0.2 400 50 $4\cdot 10^{4}$ 1000 1 1 $2.6\cdot 10^{-2}$ $3.8\cdot 10^{-4}$ $4.7\cdot 10^{-5}$ 0.991 0.983 $1.3\cdot 10^{-2}$ $7.7\cdot 10^{-3}$
@@
Table 1: Parameters and figures of merit for a circuit-QED implementation of two synchronized SPSs. We consider four parameter sets: A and B correspond state-of-the-art parameters $g,\kappa,\Delta$ and $|\alpha_{\rm ss}|^{2}$ for different values of antenna couplings $\gamma$ and decoherence rates $\gamma_{\rm loss}$, $\gamma_{\phi}$. C and D correspond to more challenging long-term parameters due to the lower coupling $g$ and larger $|\alpha_{\rm ss}|^{2}$. All parameters lead to the same effective cavity-mediated drive $\Omega_{\rm cm}/2\pi\approx 40$ MHz. The detuning of the qubits is chosen to compensate for any cavity-mediated shift, including the Lamb-shift $\delta_{q}\sim\delta_{\rm cm}$ or residual Bloch-Siegert shifts. Depending on the couplings and decoherence rates of sets A-D, the inequalities in Eq. (19) are satisfied differently, as shown in columns 8th to 10th, and this influences the performance of figures of merit in the last four columns: single-photon generation probability ${\cal P}_{1}$ [cf. Sec. III.2], synchronized two-photon generation efficiency or probability $P_{2}\approx({\cal P}_{1})^{2}$ [cf. Sec. III.3], multi-photon contamination $g^{(2)}_{\rm HBT}[0]$ [cf. Sec. III.4], and photon distinguishability $g^{(2)}_{\rm HOM}[0]$ [cf. Sec. III.5]. All these quantities are calculated using the full master equation (7), the methods in Appendices B.1 and C, and evolving up to a final time $t=15/\gamma$.
@@
Fig. 2(a)-(b) illustrates the two processes in the cavity-mediated synchronization method, using two synchronized SPSs with parameters that satisfy Eq. (19). Fig. 2(a) displays the short time dynamics: the creation of a large coherent state in the cavity which induces an oscillatory occupation around the large steady-state value, $\langle a^{\dagger}a\rangle\sim|\alpha_{\rm ss}|^{2}\sim{\color[rgb]{0,0,0}10^{2}}$ [cf. red solid line and dashed black line], while mediating the simultaneous excitation of the qubits up to $\langle\sigma_{j}^{+}\sigma_{j}^{-}\rangle\approx 1$ (when the cavity driving is switched off) [cf. blue dash-dotted line]. Fig. 2(b) shows the long-time dynamics, in which the photons are generated. Each emitter decays almost independently in a timescale $t\sim 15/\gamma$, depositing a photon into its own antenna [cf. blue dash-dotted line]. To quantify the efficiency of these emission processes, we display the single-photon and synchronized two-photon generation probabilities, ${\cal P}_{1}$ and $P_{2}$, calculated using the definitions and methods in Secs. III.2-III.3. For non-optimized parameters satisfying Eq. (19), we reach ${\cal P}_{1}\approx 0.99$ [cf. green dashed line] and $P_{2}\approx({\cal P}_{1})^{2}\approx 0.98$ [cf. yellow dotted line]. This demonstrates that both emitted photons are highly independent of each other $S_{2}\approx 1$ in Eq. (3)] and that cavity-induced correlations are indeed negligible during photon emission processes ($J_{\rm cm}/\gamma\sim{\color[rgb]{0,0,0}10^{-2}}$ and $\gamma_{\rm cm}/\gamma\sim 10^{-4}$).
@@
We performed the numerical simulations in Figs. 2(a)-(b) using both the full model in Eqs. (5)-(7) and the effective model in Eqs. (13)-(18). They show an excellent agreement with small deviations $D(x)=|x_{\rm exact}-x_{\rm eff}|<10^{-2}$ in the qubit and photon counting observables [cf. Fig. 2(c)-(d)], despite the large cavity population of $\langle a^{\dagger}a\rangle\sim{\color[rgb]{0,0,0}10^{2}}\gg 1$. This demonstrates that the strong cavity driving $\Omega_{0}\gg\kappa,|\Delta|$ does not introduce heating in the system, but only creates a large coherent state in the cavity with some fast oscillations around the mean value $|\alpha_{\rm ss}|^{2}\gg 1$ [cf. solid red and dashed black lines in Fig. 2(a)]. Since these fast oscillations have no detrimental impact on the functionality of our multiphoton generation scheme, in the remainder of the paper we use the simpler effective model (13) which uses the mean steady value (10) for describing the effect of the cavity field (see App. A for more details on the approximations involved in the model).
@@ III.1 State-of-the-art parameters
In Fig. 1(b) we sketch a possible setup to implement the synchronized multiphoton emitter using superconducting circuits. For the cavity mode $a$, we consider a superconducting transmission line resonator with decay rate $\kappa/2\pi={\color[rgb]{0,0,0}50}{\rm MHz}$, and resonance frequency $\omega_{c}$ on the order of a few GHz. These superconducting resonators support strong drives and can hold a large number of photons in steady state $|\alpha_{\rm ss}|^{2}\gg 1$ Pietikäinen et al. (2017); Jenkins et al. (2014); Macri et al. (2020), as it is required in our scheme. In particular, we consider $|\alpha_{\rm ss}|^{2}\approx 100$ of average photons in steady state, induced by a cavity driving of $\Omega_{0}/2\pi=14{\rm GHz}$ and a detuning of $\Delta/2\pi=1400{\rm MHz}$. For standard transmission line resonators such as in Ref. Pietikäinen et al. (2017), this order of driving strength $\Omega_{0}$ is obtained with a voltage source of power $P_{\rm in}\sim-50$ dBm.
@@
As two-level emitters, we consider flux-tunable transmon qubits DiCarlo et al. (2009); Barends et al. (2013); Pechal et al. (2014); Wang et al. (2020) weakly coupled to the resonator mode with strength $g/2\pi={\color[rgb]{0,0,0}4}{\rm MHz}$. The cavity mediated driving amplitude then reads $\Omega_{\rm cm}/2\pi=|\alpha_{\rm ss}|g/2\pi\approx 40$MHz. This is strong but still smaller than typical transmon anharmonicities $U/2\pi\sim 300-400{\rm MHz}$ Pechal et al. (2014); McKay et al. (2019); Zhou et al. (2020), thus preventing leakage to higher excitation states. The frequency of the qubits—which will be in the few GHz range—can be fine-tuned by sending a DC current through the antenna that is coupled to each qubit [cf. Fig. 1(b)]. This is required to bring the qubits in resonance with the external drive, and to compensate any cavity-induced shifts such as Lamb-shifts $\delta_{q}\sim\delta_{\rm cm}$ or residual Bloch-Siegert shifts Pietikäinen et al. (2017) [See Sec. V for small deviations of this condition].
The antenna of each qubit is not only used for frequency calibration, but its main purpose is to collect the photons emitted by each qubit [cf. Fig. 1(b)]. The coupling rate $\gamma$ between the qubit and antenna can be chosen to optimally fulfill the inequalities in Eq. (19) and it also determines the time-scale of photon emission $t\sim 15/\gamma$. For the circuit QED parameters discussed above, we consider two values of $\gamma$: Parameter set A with $\gamma/2\pi\approx 0.97$ MHz optimized to maximize the emission probabilities ${\cal P}_{1}\approx 0.99$ and $P_{2}\approx 0.98$, as well as parameter set B with $\gamma/2\pi=3$ MHz optimized for faster emission rate $C_{2}\approx 1.2$ MHz and lower cavity-mediated effects, $J_{\rm cm}/\gamma\sim 10^{-3}$. The value of $\gamma$ also limits the maximum unwanted decoherence rates $\gamma_{\rm loss}$ and $\gamma_{\phi}$ that can be tolerated by the device to work properly. In particular, we consider $\gamma_{\rm loss},\gamma_{\phi}\sim 2\pi\cdot 10{\rm kHz}\ll\gamma$ as realized in state-of-art experiments Barends et al. (2013); Caldwell et al. (2018); McKay et al. (2019); Andersen et al. (2020).
Table 1 summarizes all the parameter sets considered in this work and their main figures of merit as discussed below. Note that we also consider extra parameter sets C and D, which reduce cavity-mediated effects by two orders of magnitude compared to A and B, respectively. This is achieved by having a 20-fold lower cavity-emitter coupling rate $g/2\pi=0.2$ MHz compensated by a larger cavity mean photon number $|\alpha_{\rm ss}|^{2}\sim 4\cdot 10^{4}$ so that the effective drive on the qubits is preserved, $\Omega_{\rm cm}=g|\alpha_{\rm ss}|\sim 2\pi\cdot 40$MHz. Engineering such a low coupling with high homogeneity on all emitters is challenging with current circuit QED technology because the coupling becomes of the same order as parasitic capacitive couplings typically found in superconducting circuits Wang et al. (2020). Nevertheless, reducing the inhomogeneity of these parasitic couplings may be possible in the long term via highly symmetric ring resonators Huang et al. (2021); Hazra et al. (2021). Parameters C and D consider this possibility and below we discuss how the scalablity can reach up to $N\sim 1000$ SPSs in this case. The second important requirement of parameters C and D is the large mean photon displacement $|\alpha_{\rm ss}|^{2}\sim 4\cdot 10^{4}$, but this is well below the values $10^{5}-10^{9}$ photons reported with current technology Pietikäinen et al. (2017); Jenkins et al. (2014) and is achieved by having a lower detuning $\Delta/2\pi=50$ MHz, larger cavity drive $\Omega_{0}/2\pi=40$ GHz and a decay of $\kappa/2\pi=400$ MHz. We note the strong drive required by our scheme is not harmful to the superconducting device since most of its energy is reflected by the cavity and it mainly induces a coherent displacement with no thermal component (see App. A for more details). In addition, in Sec. VI.3, we show that using a long superconducting resonator $L\sim 66$cm Sundaresan et al. (2015); Chang et al. (2020) the multi-mode structure of the cavity allows to reduce the required external drive down to $\Omega_{0}/2\pi\sim 2.6$ GHz while having the same effective quantities considered in sets C and D.
In the remainder of the paper, we use the four parameter sets A-D in Table 1 to characterize the performance and scaling of the synchronized SPSs for this circuit-QED implementation.
@@
Figure 3: Few-photon statistics ${\cal P}_{n}$ and two-photon dependence $D_{2}$ as a function of various system parameters such as (a) cavity occupation $|\alpha_{\rm ss}|^{2}$, (b) qubit-cavity coupling $g/\kappa$, (c) qubit-antenna decay $\gamma/\Omega_{\rm cm}$, (d) qubit-drive detuning $(\delta-\delta_{\rm cm})/\Omega_{\rm cm}$, (e) dephasing $\gamma_{\phi}/\gamma$, and (f) relaxation loss $\gamma_{\rm loss}/\gamma$. The probabilities ${\cal P}_{0}$ (green), $1-{\cal P}_{1}$ (blue), and ${\cal P}_{2}$ (brown) are plotted using the left vertical axis, whereas the two-photon dependence $D_{2}$ (red), corresponds the right vertical axis. All quantities are computed with the effective master equation (13) using parameter set A of Table 1, except for the quantity that is changed on each subplot.